d²y/dx²+dy/dx+2y=0


1
Expert's answer
2021-04-20T03:52:51-0400

"\displaystyle
y'' + y' + 2y = 0\\
\textsf{Auxiliary equation is}\\
m^2 + m + 2 = 0\\

m = \frac{-1 \pm i\sqrt{7}}{2}\\

\therefore y = e^{-\frac{x}{2}}\left(A\cos\left(\frac{\sqrt{7}}{2}x\right) + B\sin\left(\frac{\sqrt{7}}{2}x\right)\right)"